OpenStudy (anonymous):

find x if y=-2 4x - 6/y=15 a. 9/2 b. 5 c. 3 d. -6/23

OpenStudy (anonymous):

@kah.x

OpenStudy (jdoe0001):

hint: \(\bf \cfrac{4x-6}{{\color{brown}{ y}}}=15\qquad y={\color{brown}{ -2}}\qquad \cfrac{4x-6}{{\color{brown}{ -2}}}=15\) solve for "x"

OpenStudy (anonymous):

Simply plug in -2 where you see y so your equation should look like \[\frac{ 4x-6 }{ -2 }=15\]

OpenStudy (anonymous):

it is c

OpenStudy (anonymous):